Comment Step 3 of 3 A From Table 1, the value z 1occurred in four cases ABC, ABC, BCA and ABC Calculate the logic expression for Z . Z = ABC + A BC + ΒCA + ABC =ABC + A,BC + ΒCA +(ABC + ABC + ABC) (since X+X X) =AB + BC + CA (since X +X-1) Therefore, the equation for z is AB+BC+ CA two exceptions. A pattern DELL
Answer #1

Solution:

I have explained the reason used for reduction of the boolean expression, laws that are used to reduce an expression and the solution in case of the given expression.

Reduction of a Boolean Expression ;- The complex combtnatlonal logc iscuit (ciruuta e compex wtích uut booan unchions) must de reduced wthout the change n the fun 내 on to ease the ·mplementakon. There ou some laws d to perform this redueti om A. A =0 (А) = A A·B = AtB. twe wse Hhese laws for a ven expreon to twe have the tuth toble . 0 A B C CA8C ABcAec . 1 as 0 A+A=A (Aec + A BC